# Closed-Form Interpretation of Neural Network Latent Spaces with Symbolic Gradients

### v3

@@ Abstract
It has been demonstrated in many scientific fields that artificial neural networks, like autoencoders or Siamese networks, encode meaningful concepts in their latent spaces. However, there does not exist a comprehensive framework for retrieving this information in a human-readable form without prior knowledge. In order to extract these concepts, we introduce a framework for finding closed-form interpretations of neurons in latent spaces of artificial neural networks. The interpretation framework is based on embedding trained neural networks into an equivalence class of functions that encode the same concept. We interpret these neural networks by finding an intersection between the equivalence class and human-readable equations defined by a symbolic search space. The approach is demonstrated by retrieving invariants of matrices and conserved quantities of dynamical systems from latent spaces of Siamese neural networks.
@@ 1 Introduction
The current AI revolution is driven by artificial neural networks (ANNs), particularly deep learning models. These models have enabled machines to achieve superhuman performance in a variety of tasks, such as image recognition, language translation, game playing, and even generating human-like text. However, this remarkable power comes at the expense of interpretability, often referred to as the ”black box” problem. The representational capacity of artificial neural networks relies on interactions between possibly billions of neurons. While each single neuron is easy to describe mathematically, as networks become larger, it becomes increasingly difficult to understand how these interactions give rise to a neural network’s overall prediction.
@@
Neural networks applied to perform scientific discovery are often tasked with solving problems that cannot be formulated under the umbrella of regression. Further, it is often necessary to interpret a simpler sub-concept encoded in hidden layers. For these reasons, it is desirable to have a framework capable of interpreting concepts encoded in arbitrary intermediate neurons of artificial neural networks.
@@
Prominent artificial scientific discovery methods are based on networks like autoencoders Wetzel (2017); Iten et al. (2020); Miles et al. (2021); Frohnert and van Nieuwenburg (2024) or Siamese networks Wetzel et al. (2020); Patel et al. (2022); Han et al. (2023). These networks can conceive meaningful concepts inside their latent spaces without explicit training information in the form of labeled targets. The crucial obstacle is the lack of tools that enable the recovery of such concepts without prior knowledge of them. Removing this bottleneck would hence allow scientists the discovery of potentially unknown scientific insights.
In this publication, we describe a framework that can be employed to interpret any single neuron within an artificial neural network in closed form. We demonstrate the power of our framework by rediscovering the explicit formulas of matrix invariants and conserved quantities from the latent spaces of Siamese networks.
@@
Concepts encoded in neurons in hidden layers are in general not stored in a human-readable form, but they get distorted and transformed in a highly non-linear fashion. Hence, the interpretation method is based on constructing an equivalence class around a certain neuron that contains all functions encoding the same concept as the target neuron. In practice, we interpret the neuron by searching a closed-form representative function contained in this equivalence class.
@@ 2 Related Work
The current manuscript contributes to the domain of artificial neural network interpretability with the focus of enabling new scientific discovery through latent space models. Almost all of the neural network interpretability research works in science manage to confirm or deny whether neural networks learn certain known scientific concepts. While verifying a neural network is an important task, it is unsuitable for gaining novel scientific insights. There has been limited progress toward revealing scientific insights in symbolic form from artificial neural networks that do not require previous knowledge of the underlying concept beforehand Wetzel and Scherzer (2017); Cranmer et al. (2020); Miles et al. (2021); Liu and Tegmark (2021). These cases are rare examples where the underlying concept is encoded in a linear manner or other properties of the concept simplify the interpretation problem. While there are no unified approaches to interpreting latent space models, it might in principle be possible to build such models based on architectures with symbolic layers Martius and Lampert (2016); Sahoo et al. (2018); Dugan et al. (2020); Liu et al. (2024)
Our article aims to interpret existing latent space models. We extend an interpretation framework Wetzel (2024), originally developed to interpret neural network classifiers, to interpret neural network latent spaces.
The interpretation method relies on efficiently searching the space of symbolic equations which can be achieved by genetic search algorithms which form the backend of many symbolic regression algorithms. These include Eureqa Schmidt and Lipson (2009), Operon C++ Burlacu et al. (2020), PySINDy Kaptanoglu et al. (2022), FeynBroløs et al. (2021), Gene-pool Optimal Mixing Evolutionary Algorithm Virgolin et al. (2021), GPLearn Stephens (2022) and PySR Cranmer (2023). Other symbolic regression algorithms include deep symbolic regression uses recurrent neural networks Petersen et al. (2020), symbolic regression with transformers Kamienny et al. (2022); Biggio et al. (2021) or AI Feynman Udrescu and Tegmark (2020).
@@
Figure 1: (a) Most existing approaches attempt to interpret a neural network latent space by comparing the latent with known candidate concepts. In this case, it is necessary to have the correct concept at hand, which is unsustainable for scientific discovery. (b) Our method requires only a dataset and a trained neural network to be used in conjunction with a symbolic search algorithm, which then discovers a closed-form expression describing the concept encoded in the network’s latent space.
@@ 3.1 Siamese Neural Networks
Siamese neural networks (SNN) Baldi and Chauvin (1993); Bromley et al. (1993) were originally introduced to solve fingerprint recognition and signature verification problems. SNNs consist of two identical sub-networks with shared parameters, each receiving distinct inputs which are then projected to an embedding space. These projections are then compared by a distance metric, which joins the sub-networks together at their output. Inputs belonging to the same class should obtain high similarity, while those belonging to different classes should obtain low similarity. Such a framework allows for generalization to infinite-class classification problems. The distance metric $d(\cdot)$ is chosen according to the specific problem at hand, and in our case we use the squared Euclidean distance.
@@
The anchor $x_{A}$ is the ground truth class, the positive sample $x_{P}$ is of the same class as $x_{A}$, whereas the negative sample $N$ is of a different class. Instead of using a twin network, this setup requires a triplet of identical networks, each still sharing the same weights. The triplet loss is minimized when the distance between the anchor and positive sample is minimized in the embedding space, while the distance between the anchor and negative sample is maximized. The margin parameter $\alpha$ is a positive constant which encourages separation between positive and negative samples, as $\alpha=0$ would mean that the loss could be trivially minimized by projecting all samples to the same location. Finally, the $\max(\cdot)$ operation ensures that the distance between positive and negative samples remains finite.
@@
The architecture of the sub-network $f$ depends on the underlying data. In our case, we implement $f$ as a fully-connected network.
@@ 3.2 Interpretation Framework
A neural network which is trained to perform a task requiring knowledge of some concept $g$ does not necessarily encode it in the same, human-readable format that we use. Instead, it is more likely that it learns some function $f$, which is related to $g$ by some highly nonlinear and uninterpretable transformation $\phi$,
@@
Different realizations of neural networks $f$ might learn the same concept $g$ and therefore contain the same information. More formally, these realizations are all members of the following equivalence class:
$\tilde{H}_{g}=\left\{f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})\mid\exists\text{ invertible }\phi\in C^{1}(\mathbb{R},\mathbb{R}):f(\mathbf{x})=\phi(g(\mathbf{x}))\right\}.$ (2)
While each network uses a unique invertible transformation $\phi$, they are functionally equivalent in that they learn the same concept from the data. At this point, we ask the question, whether it is possible to identify the concept $g$ without knowing the function $\phi$.
@@
In order to avoid the necessity of knowing $\phi$, we rewrite the equivalence class (2) such that membership can be defined without explicit information about $\phi$. Since all $f\in\tilde{H}_{g}$ are required to be continuously differentiable, we can show that the gradients of the two functions $f$ and $g$ point in the same direction,
@@
Here we used that $\phi$, by construction, is invertible. Since $\phi^{\prime}(g(\mathbf{x}))$ is merely a scaling factor, this equation allows us to define a new equivalence class $H_{g}\subseteq\tilde{H}_{g}$,
$H_{g}=\left\{f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})|\frac{\nabla f(\mathbf{x})}{\left\|\nabla f(\mathbf{x})\right\|}=\frac{\nabla g(\mathbf{x})}{\left\|\nabla g(\mathbf{x})\right\|}\vee\nabla f(\mathbf{x})=\nabla g(\mathbf{x})=0,\forall\mathbf{x}\in\mathcal{D}\right\}.$ (5)
In the next section, we discuss how to obtain a closed-form expression which is a member of (5).
@@ 3.3 Symbolic Search
Symbolic regression is a regression analysis technique that has traditionally been used to find closed-form expressions that approximate the relation between target and input variables for a given dataset. Typically, this is done by employing a genetic algorithm, which evolves a population of candidate formulas using genetic operations like selection, crossover, and mutation, aiming to find the least complex tree of operators $T$ that maps inputs $X$ to outputs $Y$. In the context of neural network interpretation, symbolic regression is employed to convert a complex model into an interpretable tree representation.
In our case, we search for a simple tree $T$ which represents a function $g\in H_{g}$. However, instead of performing regression on a set of prediction targets to find the best fitting function, we search for an analytical expression whose normalized gradients are as close as possible to those of $f$. Because of this difference, we refer to this approach as symbolic search instead of symbolic regression. Note that this requires that $T$ consists of operators that yield a differentiable function.
@@ 3.4 Algorithms
    Choose a neuron to interpret. This neuron computes $f_{\theta^{\prime}}(\mathbf{x})$, where $\theta^{\prime}\subseteq\theta$. Treat this as the latent model and compute its gradient with respect to the input, i.e., $\nabla_{\mathbf{x}}f_{\theta^{\prime}}(\mathbf{x})$. See algorithm 2.
@@
    Apply symbolic search to find a symbolic tree $T$ whose gradients point in the same direction as $f_{\theta}^{\prime}$. See algorithm 3.
@@
Output: Trained model $F_{\theta}$, latent model $f_{\theta^{\prime}}$
@@ 8          Backpropagate the loss and update the model parameters $\theta$
Data: Unlabelled dataset $\mathcal{D}=(X)$
Input: Trained model $F_{\theta}$, latent model $f_{\theta^{\prime}}$
@@
1 $g_{f}\leftarrow[\nabla f_{\theta^{\prime}}(\mathbf{x})\text{ for }\mathbf{x}\text{ in }X]$triangleright Evaluate gradients w.r.t. input at neuron $f$
2 $g_{f}\leftarrow[\frac{\nabla f_{\theta^{\prime}}}{\|\nabla f_{\theta^{\prime}}\|+\epsilon}\text{ for }\nabla f_{\theta^{\prime}}\text{ in }g_{f}]$triangleright Normalize Gradients

### v4

@@ Abstract
It has been demonstrated that artificial neural networks like autoencoders or Siamese networks encode meaningful concepts in their latent spaces. However, there does not exist a comprehensive framework for retrieving this information in a human-readable form without prior knowledge. In quantitative disciplines concepts are typically formulated as equations. Hence, in order to extract these concepts, we introduce a framework for finding closed-form interpretations of neurons in latent spaces of artificial neural networks. The interpretation framework is based on embedding trained neural networks into an equivalence class of functions that encode the same concept. We interpret these neural networks by finding an intersection between the equivalence class and human-readable equations defined by a symbolic search space. Computationally, this framework is based on finding a symbolic expression whose normalized gradients match the normalized gradients of a specific neuron with respect to the input variables. The effectiveness of our approach is demonstrated by retrieving invariants of matrices and conserved quantities of dynamical systems from latent spaces of Siamese neural networks.
Keywords: Artificial Neural Networks, Interpretation of Artificial Neural Networks, Symbolic Regression, Automated Scientific Discovery
@@ 1 Introduction
The current AI revolution is driven by artificial neural networks (ANNs). These models have enabled machines to achieve superhuman performance in a variety of tasks, such as image recognition, language translation, game playing, and even generating human-like text. However, this remarkable power comes at the expense of interpretability, often referred to as the black-box problem. The representational capacity of artificial neural networks relies on interactions between possibly billions of neurons. While each single neuron is easy to describe mathematically, as networks become larger, it becomes increasingly difficult to understand how these interactions give rise to a neural network’s overall prediction.
@@
Neural networks applied to perform scientific discovery are often tasked with solving problems that cannot be formulated under the umbrella of regression. Moreover, often relevant information is encoded by neurons deep inside neural networks, specifically in latent spaces. Prominent artificial scientific discovery methods have been proposed based on networks like autoencoders Wetzel (2017); Iten et al. (2020); Miles et al. (2021); Frohnert and van Nieuwenburg (2024) or Siamese networks Wetzel et al. (2020); Patel et al. (2022); Han et al. (2023). These networks can distill meaningful concepts inside their latent spaces without being supplied with explicit training information in the form of labeled targets. The crucial obstacle to their wider adoption is the lack of tools that enable the recovery of the learned concepts without prior knowledge. Removing this bottleneck would allow scientists to use these tools to discover potentially new scientific insights.
For these reasons, it is desirable to have a framework capable of interpreting concepts encoded in arbitrary intermediate neurons of artificial neural networks. In this paper, we describe such a framework that can be employed to interpret any single neuron within an artificial neural network in closed form.
The central barrier that we need to overcome is that concepts encoded in neurons in hidden layers are generally not stored in a human-readable form, but instead get distorted and transformed in a highly non-linear fashion, which limits the naive employment of symbolic regression techniques. The three obstacles towards discovering a closed-form functional interpretation of information encoded inside neural networks are:
  1. 1.
    scaling of symbolic representations: Any form of symbolic search algorithm scales poorly with the complexity of the underlying equation. Many scientists are working on competing symbolic search algorithms mainly tailored to symbolic regression, a list can be found in the subsequent paragraph.
  2. 2.
    dimensional mismatch of neural networks storing information distributed among multiple neurons. Common methods to eliminate this mismatch are based on disentangling features learned by different neurons within the same layer Higgins et al. (2017) or to enforce a bottleneck Koh et al. (2020) such that single neurons capture individual concepts.
  3. 3.
    distortions of concepts within a neural network in a highly non-linear form. If neural networks learn concepts, there is no reason to store them in a form aligned with a human formulation of the concept. For example, if a neural network learns the concept of temperature, there is no reason to choose the Celsius or the Fahrenheit scale, nor does this encoding need to be linear. In practice, it turns out that this non-linear distortion cannot even be captured with symbolic equations. This problem prevents symbolic search algorithms from interpreting anything beyond output neurons in the context of regression. Until the invention of the interpretation framework presented in our manuscript, solving this problem was impossible.
The interpretation method presented in this manuscript is currently the only option to overcome obstacle 3 and is highly complementary with other publications that overcome obstacles 1 and 2.
@@
Our proposed solution is based on the idea of retrieving information encoded inside a desired neuron instead of replicating the neuron’s output symbolically. The interpretation method is based on constructing an equivalence class around this neuron, which contains all functions encoding the same concept as the target neuron. In practice, we interpret the neuron by searching for a closed-form representative function contained in this equivalence class.
This article is structured in the following manner: After reviewing related work, we explain the mathematical basis underpinning the interpretation procedure centered around the equivalance class of all functions containing the same information as a target neuron. Further, we design an algorithm to find the simplest symbolic representative of this equivalence class. We demonstrate the power of our framework by rediscovering the explicit formulas of matrix invariants and conserved quantities from the latent spaces of Siamese networks.
@@ 2 Related Work
The current manuscript concerns the domain of artificial neural network interpretability, with a focus on enabling new scientific discovery through latent space models. Much of the neural network interpretability research adresses the question of whether or not neural networks learn certain known scientific concepts. While verifying a neural network is an important task, it is unsuitable for gaining novel scientific insights. There has been limited progress toward revealing scientific insights in symbolic form from artificial neural networks that do not require previous knowledge of the underlying concept beforehand Wetzel and Scherzer (2017); Cranmer et al. (2020); Miles et al. (2021); Liu and Tegmark (2021). These cases are rare examples where the underlying concept is encoded in a linear manner, or where other properties of the concept simplify the interpretation problem. While there are no unified approaches to interpreting latent space models, it might in principle be possible to build such models based on architectures with symbolic layers Martius and Lampert (2016); Sahoo et al. (2018); Dugan et al. (2020); Liu et al. (2024)
Our article aims instead to interpret existing latent space models. We extend an interpretation framework Wetzel (2024), originally developed to interpret neural network classifiers, to interpret neural network latent spaces.
The interpretation method relies on efficiently searching the space of symbolic equations, which can be achieved by genetic search algorithms which form the backend of many symbolic regression algorithms. These include Eureqa Schmidt and Lipson (2009), Operon C++ Burlacu et al. (2020), PySINDy Kaptanoglu et al. (2022), Feyn Broløs et al. (2021), Gene-pool Optimal Mixing Evolutionary Algorithm Virgolin et al. (2021), GPLearn Stephens (2022) and PySR Cranmer (2023). Other symbolic regression algorithms include deep symbolic regression uses recurrent neural networks Petersen et al. (2020), symbolic regression with transformers Kamienny et al. (2022); Biggio et al. (2021) or AI Feynman Udrescu and Tegmark (2020).
@@
Figure 1: (a) The Siamese network consists of two pairs of identical sub-networks $f$. From the first pair, we compute the distance between the anchor and the positive example $d(f(x_{A}),f(x_{P}))$, which should be as close to zero as possible. From the second we compute $d(f(x_{A}),f(x_{N}))$, which should be as large as possible. This facilitates a latent space where similar items are close together, while dissimlar ones are far apart. (b) Most existing approaches attempt to interpret a neural network latent space by comparing the latent with known candidate concepts. In this case, it is necessary to have the correct concept at hand, which is unsustainable for scientific discovery. (c) Our method requires only a dataset and a trained neural network to be used in conjunction with a symbolic search algorithm, which then discovers a closed-form expression describing the concept encoded in the network’s latent space.
@@ 3.1 Siamese Neural Networks
Siamese neural networks (SNN) Baldi and Chauvin (1993); Bromley et al. (1993) were originally introduced to solve fingerprint recognition and signature verification problems. SNNs consist of two identical sub-networks with shared parameters, each receiving distinct inputs which are then projected to an embedding space. These projections are then compared by a distance metric, which joins each sub-network $f$ together at their output. Inputs belonging to the same class should obtain high similarity, while those belonging to different classes should obtain low similarity. Such a framework allows for generalization to infinite-class classification problems. The distance metric $d(\cdot)$ is chosen according to the specific problem at hand, and in our case we use the squared Euclidean distance.
@@
The anchor $x_{A}$ is the ground truth class, the positive sample $x_{P}$ is of the same class as $x_{A}$, whereas the negative sample $x_{N}$ is of a different class. Instead of using a twin network, this setup requires a triplet of identical networks, each still sharing the same weights. The triplet loss is minimized when the distance between the anchor and positive sample is minimized in the embedding space, while the distance between the anchor and negative sample is maximized. The margin parameter $\alpha$ is a positive constant which encourages separation between positive and negative samples, as $\alpha=0$ would mean that the loss could be trivially minimized by projecting all samples to the same location. Finally, the $\max(\cdot)$ operation ensures that the distance between positive and negative samples remains finite.
@@
The architecture of the sub-network $f$ depends on the underlying data. In our case, we implement it as a fully-connected network. We note that our framework interprets single neurons, hence our latent layer (i.e., the final neuron in our sub-network $f$), which we wish to interpret, consists of only one neuron. The details of our architecture and training hyperparameters can be found in subsection D.2.
@@ 3.2 Interpretation Framework
The interpretation framework is designed to extract concepts in the form of symbolic equations from any single disentangled or concept bottleneck neuron within an artificial neural network. While the interpretation framework can be applied to any single neuron, for the purpose of this manuscript we perform an interpretation of the output neuron $f(\mathbf{x})$ of a single sub-net of a Siamese network defined by (3.1) which produces a scalar mapping of the input into a latent space.
$f(\mathbf{x})$ contains the full information about a certain symbolic concept $g(\mathbf{x})$ if $g(\mathbf{x})$ can be faithfully reconstructed from $f(\mathbf{x})$. Conversely, if $f(\mathbf{x})$ only contains information from $g(\mathbf{x})$ it is possible to reconstruct $f(\mathbf{x})$ from the knowledge of $g(\mathbf{x})$. In mathematical terms that means that there exists an invertible function $\phi$ such that $f(\mathbf{x})=\phi(g(\mathbf{x}))$. An example of the same concept embedded in different forms is the temperature, it can be measured in Fahrenheit or Celsius and there exists a linear transformation that maps one version of the temperature onto the other.
In general, this means that if we aim to extract information from a neural network $f$, we need to account for any nonlinear and uninterpretable transformation $\phi$ that conceals the human formulation of a concept,
@@
Different realizations of neural networks might learn the same concept $g$ and therefore contain the same information.. Each of these networks might deform the concept by a different $\phi$. More formally, these realizations are all members of the following equivalence class:
$\widetilde{H}_{g}=\left\{f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})\mid\exists\text{ invertible }\phi\in C^{1}(\mathbb{R},\mathbb{R}):f(\mathbf{x})=\phi(g(\mathbf{x}))\right\}.$ (2)
While each network $f\in\widetilde{H}_{g}$ is related to $g$ via a different unique invertible transformation $\phi$, they are functionally equivalent in that they learn the same concept from the data. At this point, we ask the question, whether it is possible to identify the concept $g$ without knowing the function $\phi$.
@@
In order to avoid the necessity of knowing $\phi$, we rewrite the equivalence class (2) such that membership can be defined without explicit information about $\phi$. Since all $f\in\widetilde{H}_{g}$ are required to be continuously differentiable, we can show that the gradients of the two functions $f$ and $g$ point in the same direction,
@@
Here we used that $\phi$, by construction, is invertible, which means $\phi$ is monotonoous. Since $\phi^{\prime}(g(\mathbf{x}))$ is merely a scaling factor, this equation allows us to define a new equivalence class $\widetilde{H}_{g}\subseteq H_{g}=H_{g+}\cup H_{g-}$, where
$H_{g\pm}=\left\{f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})|\frac{\nabla f(\mathbf{x})}{\left\|\nabla f(\mathbf{x})\right\|}=\frac{\pm\nabla g(\mathbf{x})}{\left\|\nabla g(\mathbf{x})\right\|}\vee\nabla f(\mathbf{x})=\nabla g(\mathbf{x})=0,\forall\mathbf{x}\in D\right\}.$ (5)
Trivially, if $f\in H_{g}$ then $H_{g}=H_{f}$. It can be proven that $H_{g}=\widetilde{H}_{g}$, see subsection B.1 under mild assumptions. In subsubsection B.1.1 we explore whether these assumptions are justified in typical neural network settings. In order to execute the interpretation framework we look at the definition of this equivalence class in reverse. We define an equivalence class anchored on the neural network $H_{f}$ and use a genetic algorithm to retrieve the most likely symbolic concept $g$ within $H_{f}$.
@@ 3.3 Symbolic Search
Symbolic search is primarily used in symbolic regression for finding closed-form expressions that approximate the relation between target and input variables for a given dataset. Typically, this is done by employing a genetic algorithm, which evolves a population of candidate formulas using genetic operations like selection, crossover, and mutation, aiming to find the least complex tree of operators $T$ that best maps inputs $X$ to outputs $Y$ according to some objective function. This tree consists of a set of nodes, each containing either a number, variable, or a unary or binary operator (see Figure 1 (c) for an example tree) that represent a mathematical expression. In the context of neural network interpretation, symbolic regression is employed to convert a complex model into an interpretable tree representation.
In our case, we search for a symbolic tree $T$ which represents a function $g\in H_{f+}$, meaning, we look for a symbolic concept $g$ within the equivalence class anchored on the neural network $f$. During this step we choose a symbolic quantity whose gradient points in the same direction as the gradient of the network $f$. Focussing solely on $H_{f+}$ is possible because $H_{f-}$ can be mapped to $H_{f+}$ simply by multiplying each element with 1-1. In contrast to symbolic regression, instead of performing regression on a set of prediction targets to find the best fitting function, we search for an analytical expression whose normalized gradients are as close as possible to those of $f$. Because of this difference, we refer to this approach as symbolic search instead of symbolic regression. Note that this requires that $T$ consists of operators that yield a differentiable function. To implement our symbolic search algorithm, we modify the SymbolicRegression.jl module, the backend of the PySR package Cranmer (2023).
@@
Nodes are mutated and added by the modified symbolic search algorithm in order to minimize this objective function. The unary operators we use include $\{\text{sqrt(square root)},\text{sq(square)},\text{sin},\text{exp}\}$, and for binary operators we use ${+,,,/,}\{$+,-$\wedge\}$. The setup we use is described in subsection D.1.
@@ 3.4 Algorithms
    Choose a neuron to interpret. In our specific case, we are interested in interpreting the latent space of a Siamese network $f_{\theta}$ with weights $\theta$. Compute its gradient with respect to the input, i.e., $\nabla_{\mathbf{x}}f_{\theta}(\mathbf{x})$. See algorithm 2.
@@
    Apply symbolic search to find a symbolic tree $T$ whose gradients point in the same direction as $f_{\theta}$. See algorithm 3.
@@
Output: Trained network $f_{\theta}$
@@ 8    Backpropagate the loss and update the model parameters $\theta$
Data: Unlabelled dataset $(X)$
Input: Trained network $f_{\theta}$
@@
1 $g_{f}\leftarrow[\nabla f_{\theta}(\mathbf{x})\text{ for }\mathbf{x}\text{ in }X]$ $\triangleright$ Evaluate gradients w.r.t. input at neuron $f$
[…]